### v1

# Ergodicity detection algorithms: Scaling of ergodicity in random symbolic dynamics

Abstract
The mathematical definitions of distinct concepts that are needed in building an ergodicity detection algorithm are introduced in a framework. This algorithmic framework is expressed in a discrete setting in an accessible manner for broader quantitative practitioners without loss of generality. At the same time, the common misconceptions of the requirement of visiting all available states in the time-averaged quantities for physical systems and non-existence of an ergodic process are resolved by introducing the distinction between Gibbs-Boltzmann and von Neumann-Birkhoff ergodic regimes. For this purpose, we introduce a new concept which is called sufficiency of sparse visit. We use finite symbolic random sequences as a pedagogical tool in establishing the different approaches for the detection of ergodic regimes of dynamical systems with vector patterns. The simple example system conveys the different attitudes in ergodicity regimes and offers guidance for building computational tools for its algorithmic detection.
ergodic hypothesis; randomness; symbolic dynamics;ensemble theory
pacs:
MSC:, 37A60,82B03, 82B41, 82M31, ACM:, G.3, G.2.3, I.6.1
I Introduction
The origins of the ergodic hypothesis appear to be statements put forward by Boltzmann and Maxwell, that multiple states evolving with constant energy should return to their initial conditions [1]. This simple-sounding statement as the origin of Poincaré recurrance [2] formed the basis of the connection between statistical mechanics and its thermodynamic justification, formalised by Gibbs ensemble theory [3]. However, the interpretation that a system should visit all possible states as a mapping from constant energy phases or from nearby points (quasi-ergodic) [4] to its time evolution attracted the interest of mathematical scientists [5], i.e., von Neumann and Birkhoff via their theorems [6, 7, 8], in some sense diverging from statistical physics to form a mathematical field called ergodic theory for low dimensional systems as an extension of strong form of law of large numbers [9, 10, 11]. In this context, the term ergodicity, a probe that measures if a given dynamical regime attains ensemble averages being the same as time averages or if this equivalence is broken, is studied as a key aspect of the dynamics of a physical system [12].
Measuring ergodicity is practiced and needed in many different applied quantitative fields, such as, simple liquids [13], the Ising-Lenz model with external fields and Hopfield networks [14, 15], earthquake fault dynamics [16], atmospheric forecasting [17], economics [18], financial portfolios [19], machine learning of optimal policies [20], neuromorphic computing [21], random matrices [22, 23], deep learning architecture search [24, 25], random binary sequences [26], artificial spin-ice experiments [27], relativistic geodesics [28], markov chains [29, 30, 31] and network congestion [32]. In biomedicine, sensitivity analysis in differential equations for disease trajectories would also require an attention to ergodicity [33].
In section II, we introduce the basic definition of ergodicity and distinctions we use for different kinds of ergodicity via sufficiency of sparse-visits, section III details of conceptual components of detection algorithms in the discrete setting are defined, section IV presents numerical experiments with random symbolic dynamics demonstrating ergodicity detection framework in the mentioned two different ergodicity schemes and their scaling properties. Finally, in section V, we conclude the study.
II Gibbs-Boltzmann ergodicity
A mathematical ensemble, as introduced by Gibbs differentiate itself from probabilistic notion of sample space. This was even noted in statistical inference literature [34], where ensemble entails a more generic way of building set of possibilities. In the context of statistical mechanics, this is the set of microstates, i.e., mechanical degrees of freedoms. It might be obvious that von Neumann-Birkhoff ergodicity may not be attainable for thermodynamic systems due to astronomical number of states to visit [35]. However, Gibbs-Boltzmann ergodicity states visiting most likely portion of the energy manifold in the microcanonical ensemble, this brings the concept of sufficiency of sparse visits, that makes ergodic hypothesis plausiable. Moreover, von Neumann-Birkhoff ergodicity would still hold for low-dimensional systems [36].
We define a space that a system evolves over an ensemble $\Gamma_{i}^{j}$, with different regions $i$ and formed with $j$ different states at that region. The regions are simply a collection of different states, and not overlapping. This corresponds to all possible accessible microstates as established by Gibbs. A given observable quantity $g$, as a function of ensemble members $g=g(\Gamma^{j})$, can be computed by taking its ensemble average, $\langle g\rangle^{\gamma}$, at a given sparsity level $\gamma\in(0,1)$, where 1.0 corresponds to no sparsity. In the case of sparsity level 1.0, ensemble formed with all accessible states, when some states merged and pooled regions can be formed then sparsity level is below 1.0, i.e., reduced representation of the ensemble. On the other hand, time-average of $g(t)$ is computed over states on the regions that system evolves over time on the ensemble formed at the given sparsity level. The time-averaged $g$ at the given sparsity level is expressed as,
$\langle g\rangle^{t}=\int_{t_{0}}^{t_{n}}d\Gamma_{i}^{\gamma}g(t).$
Ergodicity of the system, or approach to ergodicity at given $\epsilon$ and sparsity, is measured by the measure $\Delta$,
$\Delta_{\epsilon}=\Delta(\langle g\rangle^{\gamma},\langle g\rangle^{t})$
$\Delta$ measure could be absolute difference for example, then approach to ergodicity at a given sparsity level measured in terms of $\epsilon$,
$\epsilon^{\gamma}=|\langle g\rangle^{\gamma}-\langle g\rangle^{t}|$
Gibbs-Boltzmann ergodicity manifests at very large sparsities orders close to Avogadro’s number for macroscopic systems. This made the core of the concept of sufficiency of sparse visits, whereby very small number of visits are sufficient, compare to number of states that a system actually can attain, this is due to energy manifolds values can be constructed with very small subsets of states a system can attain [1, 13, 37]. For very low dimensional systems, the sparsity level of 1.0 corresponds to von Neumann-Birkhoff ergodicity [36]. The concept of sufficiency of sparse visits reflects the thermodynamics justification of ergodic hypothesis Boltzmann and Gibbs established with contributions from Maxwell initially [1, 3]. Coarse-graining procedure is rather distinct in representing the degrees of freedom for the system before it evolves over time, on the other hand sufficiency of sparse visits occurs over-time and it’s the property of dynamics that representation reduction occurs naturally. If a system reaches to ergodic regime in between two time points $t_{0},t_{n}]$ at the acceptable $\epsilon$, then system said to be ergodicergodic. In this sense, a process is not ergodic rather a system attains the ergodic regime at a given time interval.
Figure 1: Discrete representation of an Ensemble set with alphabet size 3 and event size 2. Symbols are visualised with black, white and gray, columns are being ensemble members.
Figure 2: Discrete representation of an Ensemble set with alphabet size 2 and event size 3. Symbols are visualised with black and white, columns are being ensemble members.
III Ergodicity detection
Discrete setting for a statistical mechanical system is not only pedagogically interesting [38], the study of dynamical system shown to be important in diverse scientific areas [39]. Because of this, we build our framework of ergodicity detection in the language of discrete dynamical systems, using the concepts introduced in the previous section.
We define a discrete dynamical system represented by the tuple for ergodicity, $\langle\mathscr{S},\mathscr{E},\mathscr{U},\mathscr{O}\rangle$, where the set of states $\mathscr{S}\in\mathbb{N}^{n}$, states can be restricted portion of $\mathscr{N}$, ensemble $\mathscr{E}$ with members $e_{j}(\mathscr{S})$, set of dynamical rules $\mathscr{U}(t)=f_{i}(e_{j})$ with $f_{i}$ being functions of ensemble members appear as recurrence equations in evolving the system from the initial states and an observable $\mathscr{O}$ as a function of ensemble members $e_{j}$. If given dynamical rules have a stochastic component then this will be random discrete dynamical system.
Figure 3: Approach to ergodicity with different alphabet and event sizes. A pattern that alphabet size and event size combined decimal number based approach to ergodicity scaling is observed.
alphabet event C $C^{se}$ $\alpha$ $\alpha^{se}$
6 5 0.4997 0.0190 -0.5059 0.0049
5 8 0.4834 0.0214 -0.4952 0.0053
5 5 0.3633 0.0189 -0.4921 0.0047
5 2 0.1765 0.0182 -0.4936 0.0046
4 8 0.3831 0.0202 -0.4946 0.0051
4 3 0.1677 0.0162 -0.4944 0.0040
3 3 0.0601 0.0207 -0.5024 0.0053
2 8 0.0429 0.0202 -0.5001 0.0052
2 5 -0.0848 0.0193 -0.4913 0.0050
2 2 -0.2310 0.0180 -0.5051 0.0046
Table 1: Scaling of approach to ergodicity for diffeernt alphabet and event sizes, Fitting the power law over $\Omega$, approach to ergodicity, $Ct^{\alpha}$, with standard error values identified with nested bootstrapping.
Figure 4: Using baseline of alphabet-event sizes 10-2, effect of deleting one symbol that corresponds to sparsity level, on the approach to ergodicity.
alphabet event sparsity C $C^{se}$ $\alpha$ $\alpha^{se}$
2 2 0.0400 -0.2555 0.0233 -0.4974 0.0054
3 2 0.0900 -0.0711 0.0219 -0.4928 0.0051
4 2 0.1600 0.1109 0.0209 -0.5021 0.0049
5 2 0.2500 0.1671 0.0208 -0.4937 0.0049
6 2 0.3600 0.2791 0.0158 -0.4988 0.0039
7 2 0.4900 0.3496 0.0213 -0.5006 0.0048
8 2 0.6400 0.4222 0.0218 -0.5008 0.0051
9 2 0.8100 0.4640 0.0219 -0.5003 0.0052
10 2 1.0000 0.5164 0.0231 -0.5014 0.0055
Table 2: Scaling of approach to ergodicity for different sparsities for the base case 10-2, the power law over $\Omega$, approach to ergodicity, $Ct^{\alpha}$, with standard error values identified with nested bootstrapping.
Ergodicity detection algorithm will have the following components,
  1. 1.
    Set up the discrete dynamical system with sparsity level $\gamma$.
  2. 2.
    Ensemble average of the observable denoted by $\langle\mathscr{O}\rangle$, computed over the ensemble $\mathscr{E}$.
  3. 3.
    Let system evolve over time with an initial conditions, in the interval $t_{0},t_{n}]$ using the dynamical rules $\mathscr{U}(t)$, and obtain set of ensemble members $\mathscr{T}(t)$, this corresponds to sampling of ensemble with replacement obeying the dynamical rules. However, $\mathscr{T}(t)$ forms a trajectory that it is an ordered set with temporal correlations,
  4. 4.
    Time average of the observable denoted by $\langle\mathscr{O}(t)\rangle$ computed over the set $\mathscr{T}(t)$. This is repeated $n$-times, in order to compute mean time average $\langle\mathscr{O}(t)\rangle_{\mu}$.
  5. 5.
    Set up an ergodicity measure $\Delta$ with $\epsilon$, deciding where system reaches to ergodic regime.
  6. 6.
    Compute ergodicity over time $\Delta(t)$, mtimesm-times and find the mean,
    $\delta(t)^{\mu}=\Delta^{\mu}(\langle\mathscr{O}\rangle,\langle\mathscr{O}(t)\rangle_{\mu})$
  7. 7.
    If $\epsilon\geq\delta(t^{e})^{\mu}$, then the system reaches to an ergodic regime at time $t^{e}$. Obviously, if the condition is still satisfied with a monotonically decreasing $\delta^{\mu}$,after $t^{e}$ till $t^{f}$, then the system is in the ergodic regime in the interval $t^{e},t^{f}]$.
If we track $\delta(t)^{\mu}$ over-time, we measure approach to ergodicity, denoted by $\Omega(t)$ for notational convenience [14]. Note that by repeating trajectories $n\cdot m$-times, we essentially applied nested bootstrap [40], which is necessary to get mathematically smooth estimate of the ergodicity measure over time, i.e., approach to ergodicity curve.
Figure 5: Using baseline of alphabet-event sizes 18-3, effect of deleting one symbol that corresponds to sparsity level, on the approach to ergodicity.
alphabet event sparsity C $C^{se}$ $\alpha$ $\alpha^{se}$
7 3 0.0588 0.4259 0.0207 -0.4971 0.0047
8 3 0.0878 0.5155 0.0211 -0.5046 0.0048
9 3 0.1250 0.5298 0.0193 -0.4955 0.0046
10 3 0.1715 0.6241 0.0227 -0.5065 0.0053
11 3 0.2282 0.6273 0.0204 -0.4977 0.0047
12 3 0.2963 0.6793 0.0205 -0.5009 0.0047
13 3 0.3767 0.7014 0.0199 -0.4965 0.0047
14 3 0.4705 0.7349 0.0203 -0.4983 0.0047
15 3 0.5787 0.7838 0.0221 -0.5023 0.0052
16 3 0.7023 0.7536 0.0208 -0.4899 0.0048
17 3 0.8424 0.7956 0.0210 -0.4935 0.0049
18 3 1.0000 0.8475 0.0215 -0.4984 0.0049
Table 3: Scaling of approach to ergodicity for different sparsities for the base case 18-3, the power law over $\Omega$, approach to ergodicity, $C\Omega^{\alpha}$, with standard error values identified with nested bootstrapping.
IV Random symbolic dynamics
Working on a discrete setting for ergodicity detection naturally bring the discrete dynamical systems and consideration of symbolic dynamics [41]. In essence, we operate on and analyse the sequences generate by an alphabet $\mathscr{A}$, which forms the states $\mathscr{S}$ of our dynamical system. We can build ensembles over symbols using different shift maps [42]. The simplest example in this context is probably a Bernoulli process, states $\mathscr{S}=\{0,1\}$. An ensemble of sequence of length 3 for these states can be represented as black and white cells. Then the ensemble looks like in Figure 2, columns being ensemble members, the set $\{000,100,010,110,001,101,011,111\}$. This can be generated via cartesian products on the alphabet repeatedly by the sequence length. Using NumPy [43], we utilize state-of-the-art random generator called PCG [44].
Similarly, we show the ensemble of $\mathscr{S}=\{0,1,2\}$ with sequence of length of 2, colour mapped white, black and gray in Figure 1.
The following components are set for random symbolic dynamics:
  1. 1.
    The tuple $\langle\mathscr{S},\mathscr{E},\mathscr{U},\mathscr{O}\rangle$ set as follows. The alphabet or states $\mathscr{S}$ are formed with the set of positive integers and zero. The ensemble set is formed via repeated cartesian product given sequence size $\mathscr{E}=\mathscr{S}\times\mathscr{S}...\times\mathscr{S}$. The function to generate trajectories is simply a random choice $\mathscr{R}$ with replacement over $\mathscr{E}$ repeatedly, $\mathscr{U}=\mathscr{R}(\mathscr{E})$. This is simplified dynamics without explicit recurrence. Observable is simply the sum of the given ensemble element $e_{j}$.
  2. 2.
    Ensemble average is computed by taking mean of the all sums of the ensemble elements. We apply given sparsity to reduce the size of ensemble if $\gamma<1.0$.
  3. 3.
    We sample the ensemble 120000 times with $\mathscr{R}$, to get a single trajectory. 40 trajectories are generated. Mean observable over trajectory is identified, i.e., sum of each ensemble member sampled.
  4. 4.
    Ergodicity measure, absolute difference between time and ensemble average is used. This is repeated 70-times.
  5. 5.
    We found the power law [45], $\Omega=Ct^{\alpha}$, for approach to ergodicity. $\alpha$ is the exponent for the scaling law and $C$ measures the strength of the approach to ergodicity. Both $\alpha$ and $C$ are computed 70 times, reporting their mean and standard errors.
For generic ergodicity detection algorithms, nested bootstrap repeats may be different.
We compute approach to ergodicity with no sparsity, corresponding to the von Neumann-Birkhoff ergodicity, for different alphabet sizes and system lengths: 65, 58, 55, 52, 48, 43, 33, 28, 25, 22. Shown in Figure 3. The rate of convergence has a smooth ordering based on alphabet size and sequence lengths. Measuring strength of the power law has similar ordering as reported in Table 1. Doubling the alphabet length approximately doubles the strenght but we see a constant scaling exponent over all lengths, about 0.5. This constant power-law scaling indicates random nature.
In order to simulate and find the effect of sparsity, i,e., Gibbs-Boltzmann ergodicity for the principle of sufficiency of sparse visits: We set a baseline systems where by we consider its full ensemble with sparsity level $\gamma=1.0$. Its sparse versions are determined by the ensemble size of smaller systems. We computed ergodicity for two baseline systems 10210-2 and 18318-3, their alphabet and sequence sizes. Results are reported on Figures 4 - 5 and Tables 2-3. We observe similar smooth variation of power-law strengths over smaller sparsity. It is shown that order of magnitude sparsity made approach to attainment of ergodicity much faster. This indicates, very large sparsities should occur in order to attain ergodic regimes for macroscopic systems.
V Conclusions
In this work we provide a clear distinction between von Neumann-Birkhoff and Gibbs-Boltzmann ergodicities reflecting the conditions where ergodic hypothesis can be attained. Along these lines, measuring and detecting ergodic regime algorithmically is formalized for the development of family of ergodicity detection algorithms. This formalisation and understanding necessary components would help practitioners in diverse fields. We have demonstrated application of the frameworks on the random symbolic dynamics, using diverse systems and identify scaling of ergodicity in these systems.
Acknowledgments: Y. Süzen for her kind support in securing research residency and encouragements on making results and tools open to research community.